OpenStudy (anonymous):

Solve the system by elimination. x + 5y - 4z = -10 2x - y + 5z = -9 2x - 10y - 5z = 0 I've been having trouble with these problems, can someone help me with this practice problem?

OpenStudy (stamp):

x + 5y - 4z = -10 2x - y + 5z = -9 2x - 10y - 5z = 0 -1(x + 5y - 4z = -10) 5(2x - y + 5z = -9) -1(2x - 10y - 5z = 0) -x - 5y + 4z = 10 10x - 5y + 25z = -45 -2x + 10y + 5z = 0 Add the equations 7x + 0y + 34z = -35 7x + 34z = -35 Do this until you can solve for x, y, or z. Then use your x, y, or z value to find the other two. It is tedious.

OpenStudy (anonymous):

Wait what did you do after the original equations?

OpenStudy (stamp):

I scaled them to obtain values that were suitable for eliminating a variable. I multiplied each part of the equation to get to the point where when I added them, the y would be eliminated and I would be left with x and z. This gets me closer to solving for a single variable than having x, y, and z.